rellie invests 12,000 pesos. A part of it earns 10% interrest annually ans the rest earns 9% interest annually. his annual income from these investments is 1, 130 pesos. how much is invested at each rate?
Answer by tutorcecilia(2152)

Use the standard form of a line.
Ax+By=c
.10x+.09(12,000-x)=1130
.10x+1080-.09x=1130 [Solve for x]
-.01x=1130-1080
x=5000
.
Y=12,000-5000=7,000
.
Check by plugging the value of x and y back into the original equation.
.10(5000)+.09(12,000-(5000))=1130
500+.09(7000)=1130
500+630=1130
1130=1130 [Checks out]